i have a couple
1. 76.26 is 23%more than what?
2. 162 is 3.6%of what?
3. Davids savingsaccount earns 2.4%APR. How long does it take for his money to double

please show me how to do these problems

Answer by stanbon!(97)

1. Let "x" be the amount you are looking for.
Then x+0.23x = 76.26
1.23x = 76.26
x = 76.26/1.23 = 62
The amount you are looking for is 62
2. Let "x" be your "what".
Then 3.6%(x)= 162
0.036x = 162
x = 162/0.036= 4500
3. Formula for simple interest is A=P(1+r)^n
A is the amount you have after some ti024me.
P is the amount you started with.
r is the annual rate of interest
n is the number of years you let the money grow.
In your case A = 2P because it doubled.
EQUATION:
2P = P(1+0.024)^n
2 = 1.024^n
Take the log or ln of both sides to get
log2 = n(log 1.024)
n = [log 2]/[log 1.024]
n = 29.23 years.
Cheers,
Stan H.